# Individual or collective treatments: how to target antimicrobial use to limit the spread of Mannheimia haemolytica among beef cattle?

### v2

@@ Introduction
Excessive and inappropriate use of antimicrobials can select resistant bacterial strains, making it harder to treat infections in both animals and humans ([woolums_multidrug_2018, chen_scoping_2022]). However, antimicrobial treatments remain the main control measure for health practioners to treat bacterial infectious diseases, thus making antimicrobioresistance a global concern ([laxminarayan_antibiotic_2013, coetzee_association_2019]).
@@
Bovine respiratory diseases (BRD) pose a significant burden on the global livestock industry, causing considerable economic losses and welfare concerns for cattle ([delabouglise_linking_2017]). BRD are complex multifactorial respiratory conditions primarily affecting young calves and feedlot cattle ([griffin_monster_2014]). They are caused by a combination of viral and bacterial pathogens, environmental factors, and stressors ([kudirkiene_occurrence_2021]). The bovine respiratory syncytial virus (BRSV) or the parainfluenza-3 (IPV-3) are frequently reported viral agents whereas common bacteria include Pasteurella multocida, Mannheimia haemolytica or Mycoplasma spp. ([caswell_mycoplasma_2010, timsit_transmission_2013, grissett_structured_2015]). They lead to reduced feed efficiency, growth rates, and overall productivity, resulting in substantial financial losses for producers ([griffin_bovine_2010, blakebrough-hall_factors_2022]). Additionally, they cause significant suffering in affected animals, characterized by symptoms such as coughing, fever, and respiratory distress ([hilton_brd_2014]). The prevention and management of BRD are crucial for maintaining animal health, welfare, and sustainable livestock production ([hilton_brd_2014]).
@@
While viral pathogens are typically more infectious, much of the internal damage and clinical signs associated with BRD arise from secondary bacterial infections ([rice_mannheimia_2007]). Pasteurellaceae such as Mannheimia haemolytica or Pasteurella multocida are often associated with severe BRD cases, based upon necropsies of animals presenting severe respiratory symptomes, as well as positive serologies ([booker_microbiological_2008, welsh_isolation_2004]). M.haemolytica is particularly recurrent, with seroprevalences up to 80% and incidence risks up to 22.7% ([assie_exposure_2009]). M. haemolytica is commonly reported during BRD episodes in the field ([klima_characterization_2012, sudaryatma_bovine_2018]). Curative treatments for BRD mostly focus on these secondary infections, as they consist in using antimicrobials ([edwards_control_2010]). M.haemolitica’s response to antimicrobials is well characterized, making it a good candidate for a pathogen specific studies investigating treatment strategies ([noyes_mannheimia_2015]).
@@
Despite considerable progress made in the livestock sector to reduce antimicrobial use (AMU) ([bateman_evaluation_1990, watts_antimicrobial_2010]) the cattle production system is still heavily reliant on this method to control BRD ([oconnor_mixed_2013, abell_mixed_2017]). Pathogen identification is not done on routine, therefore antimicrobial curative treatment is the baseline response in the event of an outbreak ([dedonder_review_2015]). Individual treatment when animals show signs of disease is mostly used to care for affected animals, but collective treatments are also widely embraced to prevent BRD outbreaks and minimize economic losses ([nickell_metaphylactic_2010, ollivett_brd_2020, baggott2011demonstration]). The latter involves the administration of appropriate medications, such as antibiotics or anti-inflammatory drugs, to affected animals as well as the entire group to prevent additional cases ([lees_rational_2002, ives_use_2015]). This solution is motivated by poor sensitivity of current BRD field diagnostic methods, typical pathogenesis of BRD, and labor issues ([ives_use_2015]). Indeed, the sensitivity for BRD diagnosis was evaluated at around 0.27, making it very unefficient to treat each animal individually upon visual appraisal ([timsit2016diagnostic]). As a result, a trade-off exists between antimicrobial use and bovine respiratory disease (BRD) occurrence. This trade-off highlights the need for responsible antimicrobial use practices, such as targeted collective treatments with antimicrobials. Yet, to this day, such strategies have not been proposed. Indeed, the relevant criteria and associated threshold to implement collective treatments are still questioned ([edwards_control_2010]).
@@
Modeling is used to better understand complex infection dynamics and compare scenarios and rank interventions ([ezanno_how_2020]). In the case of BRD, a model ranking treatment strategies at pen level has been published ([picault_modelling_2022]). It simulates the circulation of an average pathogen in conditions mimicking contrasted farming contexts such as small versus large pens (considered as isolated from other pens) and low versus high risk of contracting BRD at pen level. The authors concluded on a bigger BRD circulation in large pens composed of high risk individuals and discussed which pens could get the highest benefit from collective treatments. A second model compares the impact of farm management practices on the spread of three causal pathogens of BRD in multi-pen farms featuring several individual risk levels of contracting BRD ([sorin-dupont_modeling_2023]). This study highlighted that sorting animals in pens in accordance with their individual risk level of developing BRD participates in limiting the spread of the most contagious pathogens. However, no model integrating novel criteria for triggering collective treatments has been proposed so far.
@@
We modeled the spread of M. haemolytica in a multi-pen fattening farm and the treatment of associated BRD. We extended our stochastic and mechanistic individual-based model previously published ([sorin-dupont_modeling_2023]) for studying the circulation of specific pathogens involved in BRD, including M. haemolytica, within a fattening farm building, by including the possibility of collective treatments based on conventional and original criteria. Our model monitored 6 individual characteristics, each animal being therefore defined by its state in each of the 6 state machines included in the model. First, the model featured individual risk status of developing BRD, which referred to an individual qualitative information (low, medium, high) summarizing the level of stressors predisposing to BRD that the animal had been exposed to prior to its arrival. The model also featured hyperthermia (either due to BRD or unspecific), infection status (susceptible, asymptomatic carrier, infectious), clinical status (asymptomatic, or with mild or severe clinical signs), detection status (detected or not), and treatment status (treated or not treated). Hereafter, we detail how treatments are triggered in this new version of the model. A graphic overview of the model is proposed in Figure 1.
Figure 1: Graphic overview of the states and transitions in each process of the model. Each animal is defined by its state in each of the 6 state machines included in the model. Full arrows represent a transition from one state to another whereas the dashed lines represent regulatory links. Let $X$ be a regulator (state or transition) of transition $T_{2}$. This will be represented by a dashed arrow starting from $X$ and targeting $T_{2}$. from [sorin-dupont_modeling_2023]
@@
Similarly to previous models, animals could be detected upon exhibition of clinical signs or hyperthermia. Severe clinical signs were detected with a sensitivity of 0.6, while the sensitivity for mild clinical signs detection was averaged to 0.3, assuming that lethargy is the main driver for detection of severe clinical signs ([kayser_evaluation_2019]). We assumed a clinical check-up at every time step (12 h, in agreement with farmers’ herd monitoring frequency ([terry_strategies_2021]). After detecting the first case through visual appraisal, all hyperthermic animals were identified using rectal temperature measured at the next feeding time, 12 hours later. ([picault_modelling_2022, sorin-dupont_modeling_2023]). False discovery may occur with the temperature check as hyperthermia may be unrelated to infection.
@@
The treatment process represented a standard antimicrobial treatment protocol during the fattening period. Each animal that was detected as affected, either though visual detection or hyperthermia, received one antibiotic dose, assumed to be effective after duration $\tau_{T}$ with a probability $p_{T}$, accounting for antimicrobial resistance. Animals still exhibiting clinical signs after this duration were treated again (same duration), the number of treatments per episode being limited ($max_{T}$). Transitions from T to NT occurred in three cases: (1) the animal recovered after duration $\tau_{T}$ due to the treatment success with probability $p_{T}$; (2) the end of the infectious period occurred while under treatment but was not caused by it; (3) the animal still expressed clinical signs but was already treated $max_{T}$ times, thus was not treated again and ignored from further treatments, as per on-field recommendations ([van2017reducing]). The values of the parameters as well as their source are summarized in Table 1.
@@
Transmission rate to susceptible (/h) $\beta$ 0.005, 0.008 and 0.012 for low, medium and high risk levels respectively [picault_modelling_2022]
Spontaneous shedding probability $p_{E}$ 0.14, 0.42 and 0.68 for low, medium and high risk levels respectively [frank_colonization_1986]
Initial carrier proportion $E_{0}/N_{0}$ 0.26, 0.48 and 0.72 for low, medium and high risk levels respectively [timsit_transmission_2013]
Mild clinical sign duration $\tau_{M}$ 2-8 (med=6) days [grissett_structured_2015]
Infectious period duration $\tau_{I}$ 6-10 (med=8) days [thomas_insights_2019]
Asymptomatic period duration $\tau_{E}$ 1 day [grissett_structured_2015]
Probability of successful treatment $p_{T}$ 0.71 [dedonder_review_2015]
Probability of severe forms $p_{C}$ 0.65 [timsit_transmission_2013]
@@
The force of infection $\Phi_{i}$ for pen $i$ was frequency-dependent, taking into account the multi-pen nature of our model by accounting for the force of infection within the pen and between pens. Let $\mathbb{B}$ be the set of pens of $n$ individuals each raised simultaneously in a building. In each pen $i$, the intra pen force of infection was the sum of the individual contribution $\beta_{\rho}$ of each infectious set of individual of risk level $\rho$ ($I_{\rho,i}$). The contribution increased with the individual risk level $\rho$. Mathematically, this translates to equation (1). In a given pen $i$, the susceptible individuals experienced an intra pen force of infection with the addition of the contributions of the other pens of set $\mathbb{B}$ deprived of i ($\mathbb{B}\backslash\{i\}$) multiplied by a contact rate between pens $c$ defining how isolated the pens were from each other. In our simulations, we set $c=0.01$. This choice was motivated by the analysis of the effect of this parameter in the previous model ([sorin-dupont_modeling_2023]). This value implied a low contact rate between pens, which was compliant to the low airborne transmission of M.haemolytica ([timsit_transmission_2013]). A susceptibility factor $\sigma_{\rho}$ increasing with the individual risk level $\rho$ multiplied this total. Mathematically, this translates to equation (2). EMULSION automatically converts the force of infection $\Phi_{i}$ to a time-dependent probability $p_{\Phi_{i}}$ using the equation (3), with $\delta t$ being the current time step.
@@
$\Phi_{i}=\sigma_{\rho}\frac{\lambda_{i}+c\sum_{b\in\mathbb{B}\backslash\{i\}}{\lambda_{b}}}{N_{i}+c(N-N_{i})}$ (2)
$p_{\Phi_{i}}(\delta t)=e^{1-\Phi_{i}\delta t}$ (3)
with $\mathbb{P}$ the set of individual risk levels, $\beta_{\rho}$ the shedding level of individuals with risk level $\rho$, $I_{\rho,i}$ the number of infectious individuals with risk level $\rho$ in pen $i$, $N_{i}$ the total number of individuals in pen $i$, $N$ the total population size. This function was proposed for modeling the force of infection as it enables to represent perfectly separated pens ($c=0$) as well as a unique large pen with homogeneous contacts ($c=1$) using a single formula.
@@
As a first criterion for triggering a collective treatment (conventional criterion), we tested the cumulative incidence in detected BRD cases in the pen, a collective treatment being performed once this criterion reached a fixed threshold. Such a threshold has been conventionally used for a long time in French fattening operations and continues to be employed in this manner, with values usually around 10% of the pen size ([mornet_p_espinasse_j_veau_1977]), although these values are not backed up by dedicated field or experimental observations.
@@
To code the model, we used the open-source EMULSION framework ([picault_emulsion_2019]), which defines models in human-readable structured text files then processed by a simulation engine coded in Python. Such a modeling framework facilitates model co-development and revisions by modelers, clinicians, and epidemiologists all together. EMULSION also supports individual-based stochastic modeling, enabling simulations of small groups characterized by a high variability in biological processes. The framework allows a detailed breakdown of the modeled events and processes. Processes are articulated as finite state machines, a formalism widely employed in computer science to depict states and transitions.
@@
We investigated sixteen scenarios of BRD treatment in a building of 200 young beef animals, distributed in 10 pens of 20 animals each, during 40 days, as BRD typically occurs in the first few weeks of fattening ([timsit_transmission_2013, assie_exposure_2009]). These scenarios were defined by one of the four treatment strategies, applied in 2 possible farm-scale proportions of individual risk status of developing BRD and two ways of assigning animals to pens. The four treatment strategies were: (1) individual treatment administered upon detection of clinical signs (Individual); (2) collective treatment triggered when the pen-scale cumulative incidence of BRD cases reached the standard threshold, in addition to individual treatment (Conventional); (3) collective treatment triggered when the weighted severity reached the defined threshold, in addition to individual treatment (Severity); and (4) collective treatment triggered when the slope function reached the defined threshold, in addition to individual treatment (Slope). The two possible individual risk level distributions in the building were: (1) 30%, 40%, and 30% of low, medium, and high risk individuals respectively (HR30); and (2) 10%, 0% and 90% of low, medium and high risk individuals respectively (HR90). We chose these proportions in order to display contrasted situations, one being compliant with previously estimated risk proportion, and the other corresponding to a worst-case scenario ([amrine_evaluation_2019]). Since having a majority of low risk individuals in the building would not feature enough cases to trigger a collective treatment, it was thus left out of this study. Finally, animals could either be assigned into pens randomly (Random) or according to their individual risk level (Sorted).
@@
Each scenario had 200 stochastic replicates. At the building scale (i.e. grouping the 10 pens), we observed the distribution and the median of the cumulative incidence of BRD cases and the average time spent with severe clinical signs per affected animal (in hours) with regards to the total AMU in the building. Moreover, we compared the mortality accross the scenarios. We also computed the proportion of treatment misuse for each scenario. This measure corresponds to the ratio of false positives over predicted positives, i.e in our case the ratio between the number of doses given to healthy animals over the total AMU. Furthermore, we observed the delay between the detection of the first BRD case and the collective treatment implementation. We also observed the delay between the infection and the administration of the first dose of antimicrobial on an individual scale. Finally, for the sorted scenario, we counted the proportion of pens of each risk level that triggered a collective treatment. For each of these outputs, all applicable scenarios were compared. As we could neither assume the normality of the distributions or the homogeneity of variances across the scenarios for most of the outputs, non-parametric tests were used for the analysis. General effects of treatment modalities were assessed with Kruskal-Wallis tests and post-hoc pairwise comparisons were done by using Wilcoxon-Mann-Whitney with Bonferroni correction to reduce the family-wise error rate.
@@
To better understand the behavior of the model and to characterize the impact of parameter uncertainty, we carried out an exploration of the model sensitivity by performing a one-at-a-time sensitivity analysis on six scenarios with a common risk level proportion of 30, 40 and 30% of low risk, medium risk and high risk respectively, as it was the most realistic ([amrine_evaluation_2019]). We simulated our three collective treatment strategies in randomly allocated pens and in sorted pens.
The model was derived from a previous one on which a sensitivity analysis accounting for parameter interaction was performed ([sorin-dupont_modeling_2023]). The latter did not provide evidence that parameter interactions contributed to model output variance. Therefore, our analysis was limited to a one-by-one analysis on the thresholds for our twonew collective treatment criteria as well as the two key parameters of disease transmission: the transmission rate $\beta$ and the contact rate between pens $c$. The thresholds and the transmission rates were used at their nominal values $\pm$ 10 and 25%. We used a larger interval for the contact rate, as little effect was observed in the sensitivity analysis of the previous model: we used the nominal value (0.01), multiplied or divided by 0.1, 0.5, 5 and 10. The tested values for each parameter are summarized in Table 2.
@@
[…]

### v4

@@ Introduction
Excessive and inappropriate use of antimicrobials can select resistant bacterial strains, making it harder to treat infections in both animals and humans ([53, 11]). However, antimicrobial treatments remain the main control measure for health practioners to treat bacterial infectious diseases, thus making antimicrobioresistance a global concern ([30, 12]).
@@
Bovine respiratory diseases (BRD) pose a significant burden on the global livestock industry, causing considerable economic losses and welfare concerns for cattle ([14]). BRD are complex multifactorial respiratory conditions primarily affecting young calves and feedlot cattle ([22]). They are caused by a combination of viral and bacterial pathogens, environmental factors, and stressors ([29]). The bovine respiratory syncytial virus (BRSV) or the parainfluenza-3 (IPV-3) are frequently reported viral agents whereas common bacteria include Pasteurella multocida, Mannheimia haemolytica or Mycoplasma spp. ([10, 47, 23]). They lead to reduced feed efficiency, growth rates, and overall productivity, resulting in substantial financial losses for producers ([21, 8]). Additionally, they cause significant suffering in affected animals, characterized by symptoms such as coughing, fever, and respiratory distress ([24]). The prevention and management of BRD are crucial for maintaining animal health, welfare, and sustainable livestock production ([24]).
@@
While viral pathogens are typically more infectious, much of the internal damage and clinical signs associated with BRD arise from secondary bacterial infections ([41]). Pasteurellaceae such as Mannheimia haemolytica or Pasteurella multocida are often associated with severe BRD cases, based upon necropsies of animals presenting severe respiratory symptomes, as well as positive serologies ([9, 52]). M.haemolytica is particularly recurrent, with seroprevalences up to 80% and incidence risks up to 22.7% ([4]). M. haemolytica is commonly reported during BRD episodes in the field ([28, 43]). Curative treatments for BRD mostly focus on these secondary infections, as they consist in using antimicrobials ([15]). M.haemolitica’s response to antimicrobials is well characterized, making it a good candidate for a pathogen specific studies investigating treatment strategies ([35]).
@@
Despite considerable progress made in the livestock sector to reduce antimicrobial use (AMU) ([7, 51]) the cattle production system is still heavily reliant on this method to control BRD ([36, 1]). Pathogen identification is not done on routine, therefore antimicrobial curative treatment is the baseline response in the event of an outbreak ([13]). Individual treatment when animals show signs of disease is mostly used to care for affected animals, but collective treatments are also widely embraced to prevent BRD outbreaks and minimize economic losses ([34, 37, 5]). The latter involves the administration of appropriate medications, such as antibiotics or anti-inflammatory drugs, to affected animals as well as the entire group to prevent additional cases ([31, 25]). This solution is motivated by poor sensitivity of current BRD field diagnostic methods, typical pathogenesis of BRD, and labor issues ([25]). Indeed, the sensitivity for BRD diagnosis was evaluated at around 0.27, making it very unefficient to treat each animal individually upon visual appraisal ([46]). As a result, a trade-off exists between antimicrobial use and bovine respiratory disease (BRD) occurrence. This trade-off highlights the need for responsible antimicrobial use practices, such as targeted collective treatments with antimicrobials. Yet, to this day, such strategies have not been proposed. Indeed, the relevant criteria and associated threshold to implement collective treatments are still questioned ([15]).
@@
Modeling is used to better understand complex infection dynamics and compare scenarios and rank interventions ([16]). In the case of BRD, a model ranking treatment strategies at pen level has been published ([38]). It simulates the circulation of an average pathogen in conditions mimicking contrasted farming contexts such as small versus large pens (considered as isolated from other pens) and low versus high risk of contracting BRD at pen level. The authors concluded on a bigger BRD circulation in large pens composed of high risk individuals and discussed which pens could get the highest benefit from collective treatments. A second model compares the impact of farm management practices on the spread of three causal pathogens of BRD in multi-pen farms featuring several individual risk levels of contracting BRD ([42]). This study highlighted that sorting animals in pens in accordance with their individual risk level of developing BRD participates in limiting the spread of the most contagious pathogens. However, no model integrating novel criteria for triggering collective treatments has been proposed so far.
@@
We modeled the spread of M. haemolytica in a multi-pen fattening farm and the treatment of associated BRD. We extended our stochastic and mechanistic individual-based model previously published ([42]) for studying the circulation of specific pathogens involved in BRD, including M. haemolytica, within a fattening farm building, by including the possibility of collective treatments based on conventional and original criteria. Our model monitored 6 individual characteristics, impletemented either as input parameters or state machines. First, the model featured the individual risk status of developing BRD as an input parameter, which corresponded to an individual qualitative information (low, medium, high) summarizing the level of stressors predisposing to BRD that the animal had been exposed to prior to its arrival. The model also featured state machines defining hyperthermia (either due to BRD or unspecific), infection status (susceptible, asymptomatic carrier, infectious), clinical status (asymptomatic, or with mild or severe clinical signs), detection status (detected or not), and treatment status (treated or not treated). Hereafter, we detail how treatments are triggered in this new version of the model. A graphic overview of the model state machines is proposed in Figure 1.
Figure 1: Graphic overview of the states and transitions in each process of the model. Each animal is defined by its state in each of the 5 state machines included in the model. Full arrows represent a transition from one state to another whereas the dashed lines represent regulatory links. Let $X$ be a regulator (state or transition) of transition $T_{2}$. This will be represented by a dashed arrow starting from $X$ and targeting $T_{2}$. From [42]
@@
Similarly to previous models, animals could be detected upon exhibition of clinical signs or hyperthermia. Severe clinical signs were detected with a sensitivity of 0.6, while the sensitivity for mild clinical signs detection was averaged to 0.3, assuming that lethargy is the main driver for detection of severe clinical signs ([27]). We assumed a clinical check-up at every time step (12 h, in agreement with farmers’ herd monitoring frequency ([44]). After detecting the first case through visual appraisal, all hyperthermic animals were identified using rectal temperature measured at the next feeding time, 12 hours later. ([38, 42]). False discovery may occur with the temperature check as hyperthermia may be unrelated to infection.
@@
The treatment process represented a standard antimicrobial treatment protocol during the fattening period. Each animal that was detected as affected, either though visual detection or hyperthermia, received one antibiotic dose, assumed to be effective after duration $\tau_{T}$ with a probability $p_{T}$, accounting for antimicrobial resistance. Animals still exhibiting clinical signs after this duration were treated again (same duration), the number of treatments per episode being limited ($max_{T}$). Transitions from T to NT occurred in three cases: (1) the animal recovered after duration $\tau_{T}$ due to the treatment success with probability $p_{T}$; (2) the end of the infectious period occurred while under treatment but was not caused by it; (3) the animal still expressed clinical signs but was already treated $max_{T}$ times, thus was not treated again and ignored from further treatments, as per on-field recommendations ([49]). The values of the parameters as well as their source are summarized in Table 1.
@@
Transmission rate to susceptible for low risk infiectious individuals (/h) $\beta_{Low}$ 0.005 [38]
Transmission rate to susceptible for medium risk infectious individuals (/h) $\beta_{Medium}$ $1.5*\beta_{Low}$ [38]
Transmission rate to susceptible for high risk infectious individuals (/h) $\beta_{High}$ $1.5*\beta_{Medium}$ [38]
Spontaneous shedding probability $p_{E}$ 0.14, 0.42 and 0.68 for low, medium and high risk levels respectively [17]
Initial carrier proportion $E_{0}/N_{0}$ 0.26, 0.48 and 0.72 for low, medium and high risk levels respectively [47]
Mild clinical sign duration $\tau_{M}$ 2-8 (med=6) days [23]
Infectious period duration $\tau_{I}$ 6-10 (med=8) days [45]
Asymptomatic period duration $\tau_{E}$ 1 day [23]
Probability of successful treatment $p_{T}$ 0.71 [13]
Probability of severe forms $p_{C}$ 0.65 [47]
@@
[…]
@@
[…]
$p_{\Phi_{i}}(\delta t)=1-e^{-\Phi_{i}\delta t}$ (3)
with $\mathbb{P}$ the set of individual risk levels, $\beta_{\rho}$ the shedding level of individuals with risk level $\rho$, $I_{\rho,i}$ the number of infectious individuals with risk level $\rho$ in pen $i$, $N_{i}$ the total number of individuals in pen $i$, NN the total population size. This function was proposed for modeling the force of infection as it enables to represent perfectly separated pens ($c=0$) as well as a unique large pen with homogeneous contacts ($c=1$) using a single formula.
@@
As a first criterion for triggering a collective treatment (conventional criterion), we tested the cumulative incidence in detected BRD cases in the pen, a collective treatment being performed once this criterion reached a fixed threshold. Such a threshold has been conventionally used for a long time in French fattening operations and continues to be employed in this manner, with values usually around 10% of the pen size ([33]), although these values are not backed up by dedicated field or experimental observations.
@@
To code the model, we used the open-source EMULSION framework ([39]), which defines models in human-readable structured text files then processed by a simulation engine coded in Python. Such a modeling framework facilitates model co-development and revisions by modelers, clinicians, and epidemiologists all together. EMULSION also supports individual-based stochastic modeling, enabling simulations of small groups characterized by a high variability in biological processes. The framework allows a detailed breakdown of the modeled events and processes. Processes are articulated as finite state machines, a formalism widely employed in computer science to depict states and transitions.
@@
We investigated sixteen scenarios of BRD treatment in a building of 200 young beef animals, distributed in 10 pens of 20 animals each, during 40 days, as BRD typically occurs in the first few weeks of fattening ([47, 4]). These scenarios were defined by one of the four treatment strategies, applied in 2 possible farm-scale proportions of individual risk status of developing BRD and two ways of assigning animals to pens. The four treatment strategies were: (1) individual treatment administered upon detection of clinical signs (Individual); (2) collective treatment triggered when the pen-scale cumulative incidence of BRD cases reached the standard threshold, in addition to individual treatment (Conventional); (3) collective treatment triggered when the weighted severity reached the defined threshold, in addition to individual treatment (Severity); and (4) collective treatment triggered when the slope function reached the defined threshold, in addition to individual treatment (Slope). The two possible individual risk level distributions in the building were: (1) 30%, 40%, and 30% of low, medium, and high risk individuals respectively (HR30); and (2) 10%, 0% and 90% of low, medium and high risk individuals respectively (HR90). We chose these proportions in order to display contrasted situations, one being compliant with previously estimated risk proportion, and the other corresponding to a worst-case scenario ([3]). Since having a majority of low risk individuals in the building would not feature enough cases to trigger a collective treatment, it was thus left out of this study. Finally, animals could either be assigned into pens randomly (Random) or according to their individual risk level (Sorted).
@@
Each scenario had 200 stochastic replicates. At the building scale (i.e. grouping the 10 pens), we observed the distribution and the median of the cumulative incidence of BRD cases and the average time spent with severe clinical signs per affected animal (in hours) with regards to the total AMU in the building. Moreover, we compared the mortality accross the scenarios. We also computed the proportion of treatment misuse for each scenario. This measure corresponds to the ratio of false positives over predicted positives, i.e in our case the ratio between the number of doses given to healthy animals over the total AMU. Furthermore, we observed the delay between the detection of the first BRD case and the collective treatment implementation. We also observed the delay between the infection and the administration of the first dose of antimicrobial on an individual scale. Finally, for the sorted scenario, we counted the proportion of pens of each risk level that triggered a collective treatment. For each of these outputs, all applicable scenarios were compared. As we could neither assume the normality of the distributions or the homogeneity of variances across the scenarios for most of the outputs, non-parametric tests were used for the analysis. General effects of treatment modalities were assessed with Kruskal-Wallis tests and post-hoc pairwise comparisons were done by using Wilcoxon-Mann-Whitney with Bonferroni correction to reduce the family-wise error rate. The signficance levels are indicated throughout the figures are the following:
  • "****": p-value << 0.0001
  • "***": p-value << 0.001
  • "**": p-value << 0.01
  • "*": p-value << 0.05
  • "ns": p-value $\geq$ 0.05
@@
To better understand the behavior of the model and to characterize the impact of parameter uncertainty, we carried out an exploration of the model sensitivity by performing a one-at-a-time sensitivity analysis on six scenarios with a common risk level proportion of 30, 40 and 30% of low risk, medium risk and high risk respectively, as it was the most realistic ([3]). We simulated our three collective treatment strategies in randomly allocated pens and in sorted pens.
The model was derived from a previous one on which a sensitivity analysis accounting for parameter interaction was performed ([42]). The latter did not provide evidence that parameter interactions contributed to model output variance. Therefore, our analysis was limited to a one-by-one analysis on the thresholds for our two new collective treatment criteria as well as the two key parameters of disease transmission: the transmission rate $\beta$ and the contact rate between pens $c$. The thresholds and the transmission rates were used at their nominal values $\pm$ 10 and 25%. We used a larger interval for the contact rate, as little effect was observed in the sensitivity analysis of the previous model: we used the nominal value (0.01), multiplied or divided by 0.1, 0.5, 5 and 10. The tested values for each parameter are summarized in Table 2.
@@
Low risk transmission rate ($\beta_{Low}$) 0.005 $\{0.00375,0.0045,0.0055,0.00625\}$
Contact rate between pens ($c$) 0.01 $\{0.001,0.005,0.05,0.1\}$